OpenStudy (bbb911):

The fraction 25/100 can be reduced to 5/20, a fraction having the same value, by dividing both the numbers 25 and 100 by:

OpenStudy (pfenn1):

What would you have to divide 25 by to get 5?

OpenStudy (pfenn1):

Same thing with 100. 100/x =20 100/x=20 100=20x 100/20=x?

OpenStudy (bbb911):

u would get 5

OpenStudy (pfenn1):

right. so if you have the fraction 25/100 and you divided it by 5/5, you would get 5/20

OpenStudy (pfenn1):

You could actually reduce the fraction once more by dividing the numerator and denominator by 5. Do you know what that would be?

OpenStudy (bbb911):

0.25

OpenStudy (pfenn1):

right 5/5 = 1 20/5 = so 5/20=1/4

OpenStudy (pfenn1):

I think you have it

OpenStudy (bbb911):

so the answer should be 25 right

OpenStudy (pfenn1):

The fraction 25/100 can be reduced to 5/20, a fraction having the same value, by dividing both the numbers 25 and 100 by: The answer is 5. You would divide both 25 and 100 to get 5 and 20 25/100=5/20=1/4

OpenStudy (bbb911):

aww i see my mistake thank you very much

OpenStudy (pfenn1):

you're welcome.